# Evolution of binaries containing a hot subdwarf and a white dwarf to double white dwarfs, and double detonation supernovae with hypervelocity runaway stars

### v1

@@ 5.3 He shell masses for double white dwarf
[…]
@@ 6.1 Charateristics of double detonation supernovae
We estimated the He shell masses at ignition for systems that undergo double detonation supernova. From Sec. 4, we observe that the He masses required for double detonations range from $\sim 0.05\,\mathrm{M_{\odot}}$ to $\sim 0.33\,\mathrm{M_{\odot}}$. Previous literature suggests that the spectra from these events will be enriched with titanium (Ti), chromium (Cr), and nickel (Ni), and will not align with the typical SNe Ia spectra (Kromer et al., 2010; Fink et al., 2010; Götberg et al., 2020). According to Townsley et al. (2019), and Boos et al. (2021), thin He shells with masses around $0.01\,\mathrm{M_{\odot}}$ are required to replicate SNe Ia spectra. Therefore, these systems are unlikely to represent normal type Ia supernovae but could result in other peculiar type I supernova transients. Kupfer et al. (2022) estimated the rate of double detonation supernovae with thick He shells to be $4\times 10^{-6}\,\text{yr}^{-1}$, which is $1\%$ of the complete SNe Ia rate estimated by Sullivan et al. (2006).
@@ 6.2 Caveats and limitations
In our binary models, we assume conservative mass transfer with no mass lost from the system. This is an initial approximation, as the He retention efficiency of the accretor is still poorly constrained. Further studies are needed to understand the effects of mass loss and angular momentum loss due to mass ejection from the system. Additionally, we assume that any angular momentum gained by the accretor during the accretion process is transferred back to the orbit. These systems, when in contact, are extremely compact and are commonly expected to be tidally synchronized. Studies such as Fuller & Lai (2014) indicated that tides may be efficient enough to prevent the accretor from spinning up beyond the orbital frequency, and most of the angular momentum gained by the accretor could therefore be transferred back to the orbit. However, other studies have investigated the consequences if the WD can be significantly spun up. For example, Yoon et al. (2004) studied rotational spin-up due to He accretion in WDs and found that He shell burning is more stable in rotating WDs compared to non-rotating cases. Similarly, Neunteufel et al. (2017) investigated magnetic and rotating WDs and found that lower mass WDs ($0.54-0.8\,\mathrm{M_{\odot}}$) could accumulate 50% more mass compared to non-rotating cases before the He ignition.
In our models, stellar winds from hot subdwarfs do not affect the fate of the binary. We tested the effect of the stellar winds by implementing the wind prescription from Krtička et al. (2016).
WD explosions are a multidimensional phenomenon (Pakmor et al., 2024). Our classification between He novae and double detonation supernovae is based on a critical density criterion. However, there is a continuous distribution of ignition densities for systems undergoing double detonation supernovae and He novae, indicating that further research is necessary to understand this better. In addition, a He nova is not a final state of the system. Any system classified as He nova can further evolve to either explode the WD or become a double white dwarf. Simulating through He novae involves mass loss assumptions and is computationally expensive. Hence, more work is needed to understand this classification better. We tested the convergence of our models for the number and grid pattern of double detonation supernovae, He novae and double white dwarfs by incrementally increasing both the time step and mesh resolution.
@@ 6.3 Current and future observations
The observed system PTF J2238+743015.1 is marked in panel (b) of Fig. 6. A recent study modeling PTF J2238+743015.1 by Piersanti et al. (2024) considers the rotational spin-up of the WD due to accretion and finds that the shear heating from this process could raise the surface temperature of the accretor, making it less dense and preventing it from undergoing a double detonation supernova. Consequently, the accretor would fail to detonate the He layers during He accretion. This contrasts with the modeling of PTF J2238+743015.1 by Kupfer et al. (2022), which, using a non-rotating accretor similar to our work, proposes that the binary would end up detonating the WD. While our models do not include the spin-up of the accreting WD or possible associated shear heating, we consider it likely that tides will prevent the WD from spinning up to near critical rotation (see also arguments in Bauer & Kupfer 2021). Our models therefore allow for the possibility of a supernova detonation for this system, but the relatively low observed mass of the hot subdwarf donor favors a He nova as more likely.
@@
Another system in panel (b) of Fig. 6 is CD-30°11223 (Vennes et al., 2012; Geier et al., 2013). Modeling by Deshmukh et al. (2024) proposes that the binary would come into contact before the hot subdwarf star becomes a WD. They also predict that the accretor would undergo detonation in the He layers, which agrees with panel (b) of Fig. 6.
@@
The observed binary in panel (a) of Fig. 6 is ZTFJ2055+4651, one of the Roche-lobe filling hot subdwarf + WD binaries. Modeling by Kupfer et al. (2020b) suggests that the donor is currently in the Roche-lobe filling stage and will later evolve into a WD, resulting in a double white dwarf system. From our analysis (Fig. 6), the system could lie in the transition phase between He nova and double detonation supernova. However, since the initial orbital period is longer than the observed orbital period, it is still possible that the system would end up as a double white dwarf, in agreement with Kupfer et al. (2020b).
The other observed systems in panel (g) of Fig. 6 (Maxted et al., 2000; Geier et al., 2007; Pelisoli et al., 2021) are at the threshold of multiple possibilities, indicating that the fate of these systems is very sensitive to their initial conditions.
@@
One of the main difficulties in observing these binaries is that the companion is a massive WD. Massive WDs are difficult to observe in electromagnetic observations due to their ultra-compact nature and resulting faintness. However, a sufficiently close massive hot subdwarf is bright ($10-100\,\mathrm{L_{\odot}}$) and can be observed by photometric brightness from Gaia’s color-magnitude diagram (for the latest volume completed sample of observed hot subdwarfs, see Dawson et al. 2024). In addition, due to their short orbital periods, these systems exhibit ellipsoidal modulations due to the tidal deformation of the hot subdwarf. These can be observed using ZTF (Bellm et al., 2019), and BlackGEM (Bloemen et al., 2016), OGLE (Udalski et al., 2015) and Gaia (Gaia Collaboration et al., 2016).
Large surveys such as 4MOST (de Jong et al., 2014), WEAVE (Dalton et al., 2012), and the Milky Way mapper survey included in SDSS-V (Kollmeier et al., 2017) will provide us with RV variability to detect the binary companion.
@@ 7 Conclusion
We simulated a grid of binary models containing a hot subdwarf and a WD. Both the hot subdwarf and WD were simultaneously evolved using MESA, incorporating their binary evolution and stable mass transfer via Roche-lobe overflow. Our orbital period spans a wide period range ($36\,\text{minutes}\lesssim P_{\text{orb}}\lesssim 7.2\,\text{hours}$) to accommodate interacting and non-interacting systems. For shorter orbital periods, the binary components come into contact before the hot subdwarf becomes a WD (see Sec. 4.1 for more details). We identify several phases of mass transfer, including mass transfer during the core He burning phase, the shell He burning phase, and late thermal pulses. The initial parameter space for systems that end up as double detonation supernovae, He novae, and double white dwarfs are shown in Fig. 6. In systems where the WD explodes, the runaway velocities of the donor can reach up to $\sim\mathrm{1018\,kms^{-1}}$. Fig. 10 shows the calculated orbital velocities of donor stars as a function of the hot subdwarf mass. Systems with larger orbital periods tend to evolve into double white dwarfs. We present the most important conclusions here.
@@
    We find that the systems initiate mass transfer during core He burning, shell He burning, and thermally pulsing phase of the donor. In addition, there is also a non-negligible fraction of systems that undergo two phases of mass transfer, which leads to He ignition on the accretor. These systems begin mass transfer during the core He burning phase and later ignite the He shell of the accretor when the mass transfer rate increases during the shell-burning phase of the donor. See Sec. 4.1.1 for more details.
@@
    In addition to the runaway velocities of hot subdwarfs, we also estimate the runaway velocities of proto-WDs with a C/O core, which tend to have a higher runaway velocities compared to that of a hot subdwarf of similar mass. For more details, see Fig. 10.
@@
    We estimated the He shell masses at the time of He ignition for the systems that undergo double detonation supernova (see Fig. 9). The minimum mass required for He detonation in our models is $\sim 0.05\,\mathrm{M_{\odot}}$. In most of the cases, the ignition point is found to be above the base of the He layer.
@@
    Double white dwarf resulting from this hot subdwarf + WD binaries have thicker He shells compared to if the WDs had formed from single stars (see Fig. 11). This may affect the outcomes of double white dwarf mergers, and so Type Ia supernovae. The WD with the most massive He shell in our double white dwarf systems is a $\sim 0.88\,\mathrm{M_{\odot}}$ WD containing a $\sim 0.18\,\mathrm{M_{\odot}}$ He shell. For more details, see Sec. 4.3.1.
We compiled the existing observations of hot subdwarf + WD binaries that fall within the parameter space of our grid and marked them accordingly in Fig. 6. The marked region aligns well with the detailed modeling of these systems. Therefore, this grid will provide a first-order estimate of the potential outcomes for future hot subdwarf + WD binary observations. Since our double white dwarf properties are estimated through detailed binary evolution calculations, this grid also serves as the realistic input for 3D merger simulations, including that could lead to a Type Ia supernova.
@@ Acknowledgements.
ASR thanks Stephan Geier, Ken’ichi Nomoto, Ylva Götberg, and Jim Fuller for helpful discussions. We also express our gratitude to the Kavli Foundation for funding the MPA/Kavli Summer Program 2023, during which this collaboration became possible. Software: MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), pandas (McKinney, 2010), Grammarly, and ChatGPT (for help with grammar). Data availability: All the codes used in this project, including the MESA inlists, are available in the Zenodo and GitHub repository. The data produced during this study is available upon request.
@@ Appendix A Outliers
In this section, we present the figure explaining the evolution of properties of the two outlier systems described in section 4.3. Fig. 12 shows the evolution of properties of the binary system, comprising a $0.7\,\mathrm{M}{{}_{\odot}}$ donor, a $0.7\,\mathrm{M}{{}_{\odot}}$ accretor, and an initial orbital period of $\sim 1.3$ hours. During evolution, the accretor undergoes two episodes of accretion to accumulate $\sim 0.18\,\mathrm{M_{\odot}}$ of He. Fore more information, see Sec. 4.3.1. Fig. 13 shows the evolution of properties of a binary with initial donor mass of $0.7\,\mathrm{M}_{\odot}$, a $1.1\,\mathrm{M}_{\odot}$ accretor, and an initial orbital period of $\sim 1.7\,$hours. This system comes into contact during the later shell-burning phase of the donor and evolves into a double detonation supernova. For more details, see Sec. 4.3.2.
Figure 12: WD undergoes accretion during core He burning and shell burning phases of the donor, eventually accumulating $0.18\,\mathrm{M_{\odot}}$ of He. The panels display the evolution of different parameters as in Fig. 2. The background colors yellow, blue, and green represent the gravitational-wave inspiral phase (GW-inspiral), H mass transfer phase (H-MT), and He mass transfer phase (He-MT), respectively. $0.7\,\mathrm{M_{\odot}}$ WD has accreted $\approx 0.18\,\mathrm{M_{\odot}}$ of material producing a $\approx 0.88\,\mathrm{M_{\odot}}$ WD. For more details, see Sec. 4.3.1. This WD has the most massive He shell in our models.
Figure 13: An example of a system where a double detonation supernova occurs during the shell burning phase of the donor. The panels display the evolution of different parameters as in Fig. 2. The background colors yellow, and green represent the gravitational-wave inspiral phase (GW-inspiral), and mass transfer phase during the He shell burning phase (Shell-MT), respectively. The accretor accretes with a relatively higher accretion rate of $\sim 10^{-7}\,\mathrm{M_{\odot}\,yr^{-1}}$, resulting in ignition in the He layers less dense than the assumed critical density for detonation ($<10^{6}\,\mathrm{g\,cm^{-3}}$).

### v2

@@ 5.3 He shell masses for double white dwarf
The system that accumulates the maximum He mass ($\sim$$0.18\,\mathrm{M}_{\odot}$) without He ignition starts with a $0.7\,\mathrm{M}_{\odot}$ donor, a $0.7\,\mathrm{M}_{\odot}$ accretor, and an initial orbital period of $\sim$1.3 hours. This system undergoes two phases of mass transfer: the core He burning phase and the He shell burning phase of the donor. This system is explained in detail in Section 4.3.1. Since the He shell masses are estimated through detailed modeling of binary evolution, these systems provide realistic input for 3D merger simulations, including those that lead to Type Ia supernovae.
@@ 6.1 Charateristics of double detonation supernovae
We estimated the He shell masses at ignition for systems that undergo double detonation supernova. From Sect. 4, we observe that the He masses required for double detonations range from $\sim 0.05\,\mathrm{M_{\odot}}$ to $\sim 0.33\,\mathrm{M_{\odot}}$. Previous literature suggests that the spectra from these events will be enriched with titanium (Ti), chromium (Cr), and nickel (Ni), and will not align with the typical SNe Ia spectra (2010ApJ...719.1067K; 2010A&A...514A..53F; 2020ApJ...904...56G). According to 2019ApJ...878L..38T, and 2021ApJ...919..126B, thin He shells with masses around $0.01\,\mathrm{M_{\odot}}$ are required to replicate SNe Ia spectra. Therefore, these systems are unlikely to represent normal Type Ia supernovae but could result in other peculiar Type I supernova transients. 2022ApJ...925L..12K estimated the rate of double detonation supernovae with thick He shells to be $4\times 10^{-6}\,\text{yr}^{-1}$, which is $1\%$ of the complete SNe Ia rate estimated by 2006ApJ...648..868S.
@@ 6.2 Caveats and limitations
In our binary models, we assume conservative mass transfer with no mass lost from the system. This is an initial approximation, as the He retention efficiency of the accretor is still poorly constrained. Further studies are needed to understand the effects of mass loss and angular momentum loss due to mass ejection from the system. Additionally, we assume that any angular momentum gained by the accretor during the accretion process is transferred back to the orbit. These systems, when in contact, are extremely compact and are commonly expected to be tidally synchronized. Studies such as 2014MNRAS.444.3488F indicated that tides may be efficient enough to prevent the accretor from spinning up beyond the orbital frequency, and most of the angular momentum gained by the accretor could therefore be transferred back to the orbit. However, other studies have investigated the consequences if the WD can be significantly spun up. For example, 2004A&A...425..217Y studied rotational spin-up due to He accretion in WDs and found that He shell burning is more stable in rotating WDs compared to non-rotating cases. Similarly, 2017A&A...602A..55N investigated magnetic and rotating WDs and found that lower mass WDs ($0.54-0.8\,\mathrm{M_{\odot}}$) could accumulate 50% more mass compared to non-rotating cases before the He ignition.
We tested the effect of the stellar winds by implementing the wind prescription from 2016A&A...593A.101K. In our models, stellar winds from hot subdwarfs do not affect the fate of the binary.
Modeling thermonuclear explosions of white dwarfs is inherently a multidimensional problem. (2024A&A...686A.227P). Our classification between He novae and double detonation supernovae is based on a critical density criterion. We note that the number of systems classified as supernovae changes by less than a factor of two when the critical density threshold is changed from $3.16\times 10^{5}\ \mathrm{g\,cm^{-3}}$ to $3.16\times 10^{6}\ \mathrm{g\,cm^{-3}}$.In addition, a He nova is not a final state of the system. Any system classified as He nova can further evolve to either explode the WD or become a double white dwarf. Simulating through He novae involves mass loss assumptions and is computationally expensive. Hence, more work is needed to understand this classification better. We tested the convergence of our models for the number and grid pattern of double detonation supernovae, He novae and double white dwarfs by incrementally increasing both the time step and mesh resolution.
@@ 6.3 Current and future observations
The observed system PTF J2238+743015.1 is marked in panel (b) of Fig. 6. A recent study modeling PTF J2238+743015.1 by 2024A&A...689A.287P considers the rotational spin-up of the WD due to accretion and finds that the shear heating from this process could raise the surface temperature of the accretor, making it less dense and preventing it from undergoing a double detonation supernova. Consequently, the accretor would fail to detonate the He layers during He accretion. This contrasts with the modeling of PTF J2238+743015.1 by 2022ApJ...925L..12K, which, using a non-rotating accretor similar to our work, proposes that the binary would end up detonating the WD. While our models do not include the spin-up of the accreting WD or possible associated shear heating, we consider it likely that tides will prevent the WD from spinning up to near critical rotation (see also arguments in 2021ApJ...922..245B). Our models therefore allow for the possibility of a supernova detonation for this system, but the relatively low observed mass of the hot subdwarf donor favors a He nova as more likely.
@@
Another system in panel (b) of Fig. 6 is CD-30°11223 (2012ApJ...759L..25V; Geier2013). Modeling by 2024MNRAS.527.2072D proposes that the binary would come into contact before the hot subdwarf star becomes a WD. They also predict that the accretor would undergo detonation in the He layers, which agrees with panel (b) of Fig. 6.
@@
The observed binary in panel (a) of Fig. 6 is ZTFJ2055+4651, one of the Roche-lobe filling hot subdwarf + WD binaries. Modeling by 2020ApJ...891...45K suggests that the donor is currently in the Roche-lobe filling stage and will later evolve into a WD, resulting in a double white dwarf system. From our analysis (Fig. 6), the system could lie in the transition phase between He nova and double detonation supernova. However, since the initial orbital period is longer than the observed orbital period, it is still possible that the system would end up as a double white dwarf, in agreement with 2020ApJ...891...45K. The other observed systems in panel (g) of Fig. 6 (2000MNRAS.317L..41M; Geier2007; 2021NatAs...5.1052P) are at the threshold of multiple possibilities, indicating that the fate of these systems is very sensitive to their initial conditions.
@@
One of the main difficulties in observing these binaries is that the companion is a massive WD. Massive WDs are difficult to observe in electromagnetic observations due to their ultra-compact nature and resulting faintness. However, a sufficiently close massive hot subdwarf is bright ($10-100\,\mathrm{L_{\odot}}$) and can be observed by photometric brightness from Gaia’s color-magnitude diagram (for the latest volume completed sample of observed hot subdwarfs, see 2024A&A...686A..25D). In addition, due to their short orbital periods, these systems exhibit ellipsoidal modulations due to the tidal deformation of the hot subdwarf. These can be observed using ZTF (2019PASP..131a8002B), and BlackGEM (2016SPIE.9906E..64B), OGLE (2015AcA....65....1U) and Gaia (2016A&A...595A...1G). Large surveys such as 4MOST (2014SPIE.9147E..0MD), WEAVE (2012SPIE.8446E..0PD), and the Milky Way mapper survey included in SDSS-V (2017arXiv171103234K) will provide us with RV variability to detect the binary companion.
@@ 7 Conclusion
We simulated a grid of binary models containing a hot subdwarf and a WD. Both the hot subdwarf and WD were simultaneously evolved using MESA, incorporating their binary evolution and stable mass transfer via Roche-lobe overflow. Our orbital period spans a wide period range ($36\,\text{minutes}\lesssim P_{\text{orb}}\lesssim 7.2\,\text{hours}$) to accommodate interacting and non-interacting systems. For shorter orbital periods, the binary components come into contact before the hot subdwarf becomes a WD (see Sect. 4.1 for more details). We identify several phases of mass transfer, including mass transfer during the core He burning phase, the shell He burning phase, and late thermal pulses. The initial parameter space for systems that end up as double detonation supernovae, He novae, and double white dwarfs are shown in Fig. 6. In systems where the WD explodes, the runaway velocities of the donor can reach up to $\sim\mathrm{1018\,kms^{-1}}$. Fig. 10 shows the calculated orbital velocities of donor stars as a function of the hot subdwarf mass. Systems with larger orbital periods tend to evolve into double white dwarfs. We present the most important conclusions here.
@@
    We find that the systems initiate mass transfer during core He burning, shell He burning, and thermally pulsing phase of the donor. In addition, there is also a non-negligible fraction of systems that undergo two phases of mass transfer, which leads to He ignition on the accretor. These systems begin mass transfer during the core He burning phase and later ignite the He shell of the accretor when the mass transfer rate increases during the shell-burning phase of the donor. See Sect. 4.1 for more details.
@@
    In addition to the runaway velocities of hot subdwarfs, we also estimate the runaway velocities of proto-WDs, which tend to have a higher runaway velocities than that of a hot subdwarf of similar mass. For more details, see Fig. 10.
@@
    We estimated the He shell masses at the time of He ignition for the systems that undergo double detonation supernova (see Fig. 9). The minimum mass required for He detonation in our models is $\sim 0.05\,\mathrm{M_{\odot}}$. The white dwarf initially contains only $\sim 2\times 10^{-3}\,\mathrm{M_{\odot}}$ of He, so the majority of the He must be accreted. In most cases, the ignition point is found to be above the base of the He layer.
@@
    Double white dwarf resulting from this hot subdwarf + WD binaries have thicker He shells compared to if the WDs had formed from single stars (see Fig. 11). This may affect the outcomes of double white dwarf mergers, and so Type Ia supernovae. The WD with the most massive He shell in our double white dwarf systems is a $\sim 0.88\,\mathrm{M_{\odot}}$ WD containing a $\sim 0.18\,\mathrm{M_{\odot}}$ He shell. For more details, see Sect. 4.3.1.
We compiled the existing observations of hot subdwarf + WD binaries that fall within the parameter space of our grid and marked them accordingly in Fig. 6. The marked region aligns well with the detailed modeling of these systems. Therefore, this grid will provide a first-order estimate of the potential outcomes for future hot subdwarf + WD binary observations. Since the properties of our double white dwarf systems are derived from detailed binary evolution calculations, this grid can serve as an input for future 3D merger simulations that aim to model potential Type Ia supernova outcomes.
@@ Acknowledgements.
ASR thanks Stephan Geier, Ken’ichi Nomoto, Ylva Götberg, and Jim Fuller for helpful discussions. We also express our gratitude to the Kavli Foundation for funding the MPA Kavli Summer Program 2023, during which this collaboration became possible. ASR thanks Simon Jeffrey for bringing Ton S 415 to our attention. Software: MESA (Paxton2011; Paxton2013; Paxton2015; 2018ApJS..234...34P; Paxton2019; 2023ApJS..265...15J), Matplotlib (2007CSE.....9...90H), NumPy (2011CSE....13b..22V), pandas (mckinney2010). Data availability: All the codes used in this project, including the MESA inlists, are available in the Zenodo and GitHub repository. The data produced during this study is available upon request.
@@ Appendix A Outliers
In this section, we present the figure explaining the evolution of properties of the two outlier systems described in section 4.3. Fig. 12 shows the evolution of properties of the binary system, comprising a $0.7\,\mathrm{M}{{}_{\odot}}$ donor, a $0.7\,\mathrm{M}{{}_{\odot}}$ accretor, and an initial orbital period of $\sim 1.3$ hours. During evolution, the accretor undergoes two episodes of accretion to accumulate $\sim 0.18\,\mathrm{M_{\odot}}$ of He. For more information, see Sect. 4.3.1. Fig. 13 shows the evolution of properties of a binary with initial donor mass of $0.7\,\mathrm{M}_{\odot}$, a $1.1\,\mathrm{M}_{\odot}$ accretor, and an initial orbital period of $\sim 1.7\,$hours. This system comes into contact during the later shell-burning phase of the donor and evolves into a double detonation supernova. For more details, see Sect. 4.3.2.
Figure 12: WD undergoes accretion during core He burning and shell burning phases of the donor, eventually accumulating $0.18\,\mathrm{M_{\odot}}$ of He. The panels display the evolution of different parameters as in Fig. 2. The background colors yellow, blue, and green represent the gravitational-wave inspiral phase (GW-inspiral), H mass transfer phase (H-MT), and He mass transfer phase (He-MT), respectively. $0.7\,\mathrm{M_{\odot}}$ WD has accreted $\approx 0.18\,\mathrm{M_{\odot}}$ of material producing a $\approx 0.88\,\mathrm{M_{\odot}}$ WD. For more details, see Sect. 4.3.1. This WD has the most massive He shell in our models.
Figure 13: Example of a system where a double detonation supernova occurs during the shell burning phase of the donor. The panels display the evolution of different parameters as in Fig. 2. The background colors yellow, and green represent the gravitational-wave inspiral phase (GW-inspiral), and mass transfer phase during the He shell burning phase (Shell-MT), respectively. This system begins interacting during a later expansion phase, where the donor’s expansion rate is slower, resulting in lower accretion rates of around $10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$. This lower accretion rate leads to He ignition in the dense ($10^{6}\,\mathrm{g\,cm^{-3}}$) layers of the WD, ultimately getting classified as a double detonation supernova. For more details, see Sect. 4.3.2.